OpenStudy (anonymous):

Use the substitution method to solve the system of equations. 2x – 5y = –14 x – 2y = –5

OpenStudy (anonymous):

x=3 y=4

OpenStudy (anonymous):

Equation 1: 2x-5y = -14 Equation 2: x-2y = -5 You can solve for x or y using either equation 1 or 2. I'll solve for x using Equation 2: Step 1: Solve for x (means you want x=something) Equation 2: x-2y= -5 x=-5+2y (Added 2y to both left and right hand sides) Step 2: Substitute x=-5+2y into Equation 1 and solve for y Equation 1: 2x-5y =-14 2(-5+2y) -5y=-14 substitute x=-5+2y -10+4y-5y=-14 Distribute the 2 -10-y=-14 Simplify -y=-14+10 Added 10 to both sides -y=-4 Simplify y=4 Divide both sides by -1 Step 3: Using y=4 plug this back into Equation 1 or Equation 2 and solve for x (now you'll get x= a number) Equation 2: x-2y=-5 x-2(4)=-5 Substitute y=4 x-8=-5 simplify x=-5+8 added 8 to both sides x=3 simplify So your solution is x=3 and y=4 or another way to state this is (3,4) is a solution of the system of equations. Another way to say this is, the lines given by Equation 1 and Equation 2 intersect at the point (3,4)